OpenStudy (anonymous):

here is the next one. Use the appropriate form of the percentage formula. What percent of 8 is 2?

OpenStudy (amistre64):

what is the percentage formula?

OpenStudy (amistre64):

Part --------- ; is a good set up to remember % | Total

OpenStudy (amistre64):

% = part/total total = part/% Part = % times total

OpenStudy (amistre64):

since they give you a part and a total in the question, and ask you to find the "other", id use % = pert/total

OpenStudy (amistre64):

....part/total even :)

OpenStudy (anonymous):

Here's another way of looking at this: Translate this question from English to Math: Let x = What percentage? "of" means to multiply What percent of 8 is 2? x * 8 = 2 x = 2/8 x = 1/4 x = .25 x = 25%